# Thompson sampling MAB multi-armed bandit

### v1

@@ Abstract
This paper focuses on selecting the arm with the highest variance from a set of $K$ independent arms. Specifically, we focus on two settings: (i) regret setting, that penalizes the number of pulls of suboptimal arms in terms of variance, and (ii) fixed-budget best arm identification (BAI) setting, that evaluates the ability of an algorithm to determine the arm with the highest variance after a fixed number of pulls. We develop a novel online algorithm called UCB-VV for the regret setting and show that its upper bound on regret for bounded rewards evolves as $\mathcal{O}\left(\log{n}\right)$ where $n$ is the horizon. By deriving the lower bound on the regret, we show that UCB-VV is order optimal. For the fixed budget BAI setting and propose the SHVV algorithm. We show that the upper bound of the error probability of SHVV evolves as $\exp\left(-\frac{n}{\log(K)H}\right)$, where $H$ represents the complexity of the problem, and this rate matches the corresponding lower bound. We extend the framework from bounded distributions to sub-Gaussian distributions using a novel concentration inequality on the sample variance. Leveraging the same, we derive a concentration inequality for the empirical Sharpe ratio (SR) for sub-Gaussian distributions, which was previously unknown in the literature. Empirical simulations show that UCB-VV consistently outperforms $\epsilon$-greedy across different sub-optimality gaps though it is surpassed by VTS, which exhibits the lowest regret, albeit lacking in theoretical guarantees. We also illustrate the superior performance of SHVV, for a fixed budget setting under 6 different setups against uniform sampling. Finally, we conduct a case study to empirically evaluate the performance of the UCB-VV and SHVV in call option trading on $100$ stocks generated using geometric Brownian motion (GBM).
@@ I Introduction and Motivation
For portfolio managers, maximizing the frequency of selecting the highest variance option unlocks opportunities to optimize returns while leveraging the asymmetrical payoff structure of call options from favorable price movements. For example, for a call option buyer, a higher variance is particularly advantageous as it increases the potential for substantial gains while capping the maximum loss to the premium paid for the option. In this paper, we investigate this aspect under the regret setting by developing a bandit algorithm UCB-VV for call option trading using the variance associated with a stock. It must be noted that the results of our framework can be extended to the setting to pick the arm with the minimum variance as well.
@@ I-A2 Literature in finance
In the context of finance, variance is used to model risk premium, providing insights into market sentiment and investor behavior. Specifically, the variance risk premium (a compensation that investors demand for bearing the risk associated with the volatility in the returns of an asset) plays a critical role in both theoretical and practical applications. This concept has been leveraged to develop trading strategies, such as selling volatility when the premium is high, as demonstrated in [27, 28]. Beyond its application in trading strategies, variance is also instrumental in understanding cross-market dynamics. For instance, Hamao et al. [29] investigate how volatility in one market, such as Japan, influences volatility in another, such as the U.S., using variance as a metric to assess volatility spillovers. Similarly, Engle et. al [30] examine intraday volatility in foreign exchange markets, employing variance to model the clustering of volatility over time. This clustering, a manifestation of the heteroscedastic behavior, is effectively captured by variance. Variance has been central to research on asset pricing and portfolio management. French et al. [31] explore the relationship between expected stock returns and volatility, with volatility quantified by variance, thereby informing optimal portfolio construction and asset pricing. Collectively, these studies illustrate the multifaceted role of variance in financial research and practice.
@@ II Problem formulation for regret
Consider a MAB framework with $K\geq 2$ arms, a single player, and a time horizon $n$. We assume that the underlying rewards distributions of the arms are $\nu_{1},\nu_{2},\dots,\nu_{K}$, each of which is bounded [^1] In Section IV we relax the boundedness assumption and demonstrate the extension of our framework to sub-Gaussian distributions. in $[l,u]$. The means, $\{\mu_{i}\}$, and the variances, $\{\sigma_{i}^{2}\}$, of the distributions are fixed but unknown to the player. When playing an arm $i$ at time $t$, a random reward $X_{i}(t)\sim\nu_{i}$. The reward samples are independent across arms and time slots. A policy $\pi$ orchestrates the action of the player at each time instant and is driven by the historical data. Let $s_{i}(n)$ represent the number of plays of an arm $i$ until (and including) time $n$ by the player. In a standard stochastic bandit problem, the regret setting focuses on optimizing the trade-off between exploration and exploitation to maximize the expected cumulative reward over time. Without loss of generality, we assume that the optimal arm is unique and $\sigma^{2}_{\ast}=\max_{i\in[K]}\sigma_{i}^{2}$.
@@
The biased empirical variance of an arm $i$ at time $t$, calculated from the i.i.d. samples $\{X_{i}(s_{i}(q))\}_{q=1}^{t}$ drawn from its distribution $\nu_{i}$ given as
@@
The regret defined in (2) deviates from the classical definition of regret (e.g., see [33]) that considers the regret as the expected deviation of the metric obtained by the algorithm from its optimal value. Unlike the mean, for maximizing the variance, the optimal policy may not be equivalent to sampling the arm with the highest variance at all times, but consists of strategic switching between the arms [34]. However, since our focus is on selection of the arm with the highest variance, and penalizing a policy each time it selects a variance sub-optimal arm, (2) is an appropriate choice of measuring the performance.
@@ II-A UCB-VV Algorithm
comprising of two distinct components: $\bar{V}_{i}(t)$, represents the estimated variance, while the second term, ${\sqrt{\frac{2\log{t}}{s_{i}(t)}}=\left(C_{t,s_{i}}\right)}$, corresponds to its associated confidence bound.
@@ II-B Upper bound on regret of UCB-VV
$\mathcal{R}_{VV}(n)\leq 8\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\frac{\log{n}}{\delta_{i}}+\left(1+\frac{\pi^{2}}{3}\right)\left(\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\delta_{i}\right).$
@@
Algorithm 2 Call execution with UCB-VV
Input: Output of Algorithm 1 : Top 10 arms
Output:Rewards
1:  for each $t=0,1,\dots,N$ do
2:     for each i = 1: Top 10 do
3:        if $S_{t}>R$ then
4:           Play arm i
5:           Collect reward.
6:        else
7:           Hold
8:        end if
9:     end for
10:  end for
@@ II-C Lower bound on regret
In this section, we establish the order optimality of the proposed algorithm UCB-VV by deriving the lower bound of the problem, and show it to be matching up to constant or logarithmic factors with the upper bound of $\mathcal{O}\log(n)$. For that, we start with a couple of basic definitions.
@@
A policy $\pi$ is $\alpha$-consistent $10<\alpha<$ 1 if for one parameter distributions and for all $i\neq*$, $\mathbb{E}[s_{i}(n)]\leq n^{\alpha},$ where * represents the optimal arm.
@@
Consider the MAB problem where the objective is to optimize the variance of the observed rewards. Let $\pi$ be an $\alpha$-consistent policy and $\Theta=\left\{\left\{\theta_{i}\right\}_{i=1}^{K}\colon\theta_{i}\in\mathcal{U}\right\}$ be the distribution model. Under Assumption 1, the model-specific regret satisfies, for any constant $C_{1}<1-\alpha$, $10<C_{2}<$ 1, with $f_{i}$ and $f_{*}$ as the reward distribution of the i-th arm and the optimal arm respectively, we have
$\liminf_{n\rightarrow\infty}\frac{R(n)}{\log n}\geq\sum_{i=1,i\neq*}^{K}\frac{C_{1}C_{2}\delta_{i}}{I(f_{i},f_{*})}.$
@@ III Problem Formulation for BAI
Next, we examine the challenge of BAI within a fixed budget. The goal of the learner is to identify and recommend the arm with the highest variance using $n$ samples. Specifically, we propose an algorithm called SHVV, that draws motivation from the foundational framework of sequential halving algorithm introduced by Karnin et al [17]. Specifically, we explore a scenario where we have a predetermined budget of $n$, aiming to maximize the likelihood of accurate identification of the highest variance arm. The total budget $n$ is divided equally into $\log_{2}(K)$ elimination rounds. Let $A_{r}$ be the set of arms in round $r$, which are pulled an equal number of times. After completing a round, we eliminate half of the arms with the lowest estimated variance. The goal of BAI setting is to minimize the error probability, i.e., recommending a sub-optimal arm at the end of the time horizon, given as,
@@
The SHVV algorithm is given in Algorithm 3.
@@ III-A Upper bound on error for BAI
$\mathbb{P}(\bar{V}_{i}(t_{r})>\bar{V}_{*}(t_{r}))=\mathbb{P}\bigg{(}\frac{1}{t_{r}-1}\sum_{q=1}^{t_{r}}(X_{i}(q)-\bar{X}_{i})^{2}-$
$\frac{1}{t_{r}-1}\sum_{q=1}^{t_{r}}(X_{*}(q)-\bar{X}_{*})^{2}-(\sigma^{2}_{i}-\sigma^{2}_{*})>\delta_{i}\bigg{)}$
@@
$e_{n}\leq 3\log_{2}(K)\exp\left(\frac{-(n-\log_{2}(K)^{2})}{8n\log_{2}(K)u^{4}H_{2}}\right)$
@@
$3\sum_{r=1}^{\log_{2}(K)}\exp\left(-\frac{(n-4i_{r}\log_{2}(K))^{2}\delta_{i_{r}}^{2}}{8ni_{r}\log_{2}(K)}\right)$
@@
$\leq 3\log_{2}(K)\exp\left(\frac{-(n-K\log_{2}(K))^{2}}{8n\log_{2}(K)H_{2}}\right).$
@@ III-B Lower bound on error for BAI
$\mathbb{P}\left(\left|\frac{1}{t}\sum Z_{s}-\mathbb{E}[Z]\geq\epsilon\right|\right)\leq 2\exp\left(\frac{-2{n}\epsilon^{2}}{M^{2}}\right).$
@@
$\hskip 56.9055pt\frac{1}{6}\exp\left(-60\frac{n}{H_{\sigma}(1)\sigma_{\nu}^{2}}-\frac{M}{\sqrt{2}}\sqrt{n\log(6nK)}\right)$
where $\mathbb{P}_{k}(J_{n}\neq k)$ is the probability of error under problem $G^{k}$, $H(k)=\sum_{1\leq i\leq K,i\neq k}(\delta_{i})^{-2}$ measures the difficulty of identifying the best arm in problem $G^{k}$, $H_{\sigma}(1)=H(1)\delta_{k}$.
@@ IV Sub-Gaussian Distributions and Sharpe Ratio
In this section, we extend our framework from bounded to sub-Gaussian distributions. In particular, we derive a concentration inequality for the sample variance of sub-Gaussian random variables, followed by extending the analysis to establish novel concentration bounds for the SR.
@@
Figure 1: Regret v/s time for UCB-VV, $\epsilon$-Greedy and VTS for $\delta=0.1$ and (b) $\delta=0.5$ (c) UCB-VV v/s KL-UCB.
@@ IV-B UCB-Sharpe algorithm and its regret
[…]
@@
6:        Set index: $\text{Index}_{i}(t)=\bar{S}_{i}(t)+\beta_{i}(t)$
@@
$\text{R}_{\rm SR}(n)=\sum_{i:\Delta_{i}>0}\Delta_{i}\cdot\mathbb{E}[s_{i}(n)],$ (4)
@@
$\text{R}_{\rm SR}(n)\leq\sum_{i:\Delta_{i}>0}\left(\frac{9}{c\Delta_{i}}\log n+\mathcal{O}(1)\right).$
@@
The regret scales logarithmically with $n$, similar to classical UCB algorithms, with constants depending on the sub-Gaussian parameter $v^{2}$ and sub-optimality gap $\Delta_{i}$.
Figure 2: (a) Error probability $e_{n}$ of SHVV and uniform sampling algorithm for 6 experimental setups defined.

### v2

@@ Abstract
This paper focuses on selecting the arm with the highest variance from a set of $K$ independent arms. Specifically, we focus on two settings: (i) regret setting, that penalizes the number of pulls of suboptimal arms in terms of variance, and (ii) fixed-budget best arm identification (BAI) setting, that evaluates the ability of an algorithm to determine the arm with the highest variance after a fixed number of pulls. We develop a novel online algorithm called UCB-VV for the regret setting and show that its upper bound on regret for bounded rewards evolves as $\mathcal{O}\left(\log{n}\right)$ where $n$ is the horizon. By deriving the lower bound on the regret, we show that UCB-VV is order optimal. For the fixed budget BAI setting we propose the SHVV algorithm. We show that the upper bound of the error probability of SHVV evolves as $\exp\left(-\frac{n}{\log(K)H}\right)$, where $H$ represents the complexity of the problem, and this rate matches the corresponding lower bound. We extend the framework from bounded distributions to sub-Gaussian distributions using a novel concentration inequality on the sample variance. Leveraging the same, we derive a concentration inequality for the empirical Sharpe ratio (SR) for sub-Gaussian distributions, which was previously unknown in the literature. Empirical simulations show that UCB-VV consistently outperforms $\epsilon$-greedy across different sub-optimality gaps though it is surpassed by VTS, which exhibits the lowest regret, albeit lacking in theoretical guarantees. We also illustrate the superior performance of SHVV, for a fixed budget setting under 6 different setups against uniform sampling. Finally, we conduct a case study to empirically evaluate the performance of the UCB-VV and SHVV in call option trading on $100$ stocks generated using geometric Brownian motion (GBM).
@@ I Introduction and Motivation
For portfolio managers, maximizing the frequency of selecting the highest variance option unlocks opportunities to optimize returns while leveraging the asymmetrical payoff structure of call options from favorable price movements. For example, for a call option buyer, a higher variance is particularly advantageous as it increases the potential for substantial gains while capping the maximum loss to the premium paid for the option. In this paper, we investigate this aspect by developing bandit algorithms for call option trading using the variance associated with a stock. It must be noted that the results of our framework can be extended to the setting to pick the arm with the minimum variance as well.
@@ I-A2 Literature in finance
In the context of finance, variance is used to model risk premium, providing insights into market sentiment and investor behavior. Specifically, the variance risk premium (a compensation that investors demand for bearing the risk associated with the volatility in the returns of an asset) plays a critical role in both theoretical and practical applications. This concept has been leveraged to develop trading strategies, such as selling volatility when the premium is high, as demonstrated in [27, 28]. Beyond its application in trading strategies, variance is also instrumental in understanding cross-market dynamics. For instance, Hamao et al. [29] investigated how volatility in one market, such as Japan, influences volatility in another, such as the U.S., using variance as a metric to assess volatility spillovers. Similarly, Engle et. al [30] examined intraday volatility in foreign exchange markets, employing variance to model the clustering of volatility over time. This clustering, a manifestation of the heteroscedastic behavior, is effectively captured by variance. Variance has been central to research on asset pricing and portfolio management. French et al. [31] explored the relationship between expected stock returns and volatility, with volatility quantified by variance, thereby informing optimal portfolio construction and asset pricing. Collectively, these studies illustrate the multifaceted role of variance in financial research and practice.
@@ II Problem formulation for regret
Consider a MAB framework with $K\geq 2$ arms, a single player, and a time horizon $n$. We assume that the underlying rewards distributions of the arms are $\nu_{1},\nu_{2},\dots,\nu_{K}$, each of which is bounded [^1] In Section IV, we relax the boundedness assumption and demonstrate the extension of our framework to sub-Gaussian distributions. in $[l,u]$. The means, $\{\mu_{i}\}$, and the variances, $\{\sigma_{i}^{2}\}$, of the distributions are fixed but unknown to the player. When playing an arm $i$ at time $t$, a random reward $X_{i}(t)\sim\nu_{i}$ is received. The reward samples are independent across arms and time slots. A policy $\pi$ orchestrates the action of the player at each time instant and is driven by the historical data. Let $s_{i}(n)$ represent the number of plays of an arm $i$ until (and including) time $n$ by the player. In a standard stochastic bandit problem, the regret setting focuses on optimizing the trade-off between exploration and exploitation to maximize the expected cumulative reward over time. Without loss of generality, we assume that the optimal arm is unique and $\sigma^{2}_{\ast}=\max_{i\in[K]}\sigma_{i}^{2}$.
@@
The biased empirical variance of an arm $i$ at time $t$, calculated from the i.i.d. samples $\{X_{i}(s_{i}(q))\}_{q=1}^{t}$ drawn from its distribution $\nu_{i}$ is given as
@@
The regret defined in (2) deviates from the classical definition of regret (e.g., see [33]) that considers the regret as the expected deviation of the metric obtained by the algorithm from its optimal value. Unlike the mean, for maximizing the variance, the optimal policy may not be equivalent to sampling the arm with the highest variance at all times, but consists of strategic switching between the arms [34]. However, since our focus is the selection of the arm with the highest variance, and penalizing a policy each time it selects a variance sub-optimal arm, (2) is an appropriate choice of measuring the performance.
@@ II-A UCB-VV Algorithm
comprises of two distinct components: $\bar{V}_{i}(t)$, represents the estimated variance, while the second term, $C_{t,s_{i}}\triangleq{\sqrt{\frac{2\log{t}}{s_{i}(t)}}}$, corresponds to its associated confidence bound.
@@ II-B Upper bound on regret of UCB-VV
$\text{R}_{\texttt{VV}}(n)\leq 8\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\frac{\log{n}}{\delta_{i}}+\left(1+\frac{\pi^{2}}{3}\right)\left(\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\delta_{i}\right).$
@@ II-C Lower bound on regret
In this section, we establish the order optimality of the proposed algorithm UCB-VV by deriving the lower bound of the problem, and show it to be matching up to constant factors with the upper bound of $\mathcal{O}\log(n)$. For that, we start with a couple of basic definitions.
@@
A policy $\pi$ is $\alpha$-consistent $10<\alpha<$ 1, if for one parameter distributions and for all $i\neq*$, $\mathbb{E}[s_{i}(n)]\leq n^{\alpha},$ where * represents the optimal arm.
@@
Consider the MAB problem where the objective is to optimize the variance of the observed rewards. Let $\pi$ be an $\alpha$-consistent policy and $\Theta=\left\{\left\{\theta_{i}\right\}_{i=1}^{K}\colon\theta_{i}\in\mathcal{U}\right\}$ be the distribution model. Under Assumption 1, for any constant $C_{1}<1-\alpha$, $10<C_{2}<$ 1, with $f_{i}$ and $f_{*}$ as the reward distribution of the i-th arm and the optimal arm respectively, the model-specific regret satisfies,
$\liminf_{n\rightarrow\infty}\frac{\text{R}_{\texttt{VV}}(n)}{\log n}\geq\sum_{i=1,i\neq*}^{K}\frac{C_{1}C_{2}\delta_{i}}{I(f_{i},f_{*})}.$
@@ III Problem Formulation for BAI
Next, we examine the challenge of BAI within a fixed budget. The goal of the learner is to identify and recommend the arm with the highest variance using $n$ samples. Specifically, we propose an algorithm called SHVV, that draws motivation from the foundational framework of sequential halving algorithm introduced by Karnin et al [17]. We explore a scenario where we have a predetermined budget of $n$, aiming to maximize the likelihood of accurate identification of the highest variance arm. The total budget $n$ is divided equally into $\log_{2}(K)$ elimination rounds. Let $A_{r}$ be the set of arms in round $r$, which are pulled an equal number of times. After completing a round, we eliminate half of the arms with the lowest estimated variance. The goal of BAI setting is to minimize the error probability, i.e., recommending a sub-optimal arm at the end of the time horizon, given as,
@@
Here $J_{n}$ represents the recommendation. The SHVV algorithm is given in Algorithm 2.
@@ III-A Upper bound on error for BAI
$\mathbb{P}(\bar{V}_{i}(t_{r})>\bar{V}_{*}(t_{r}))=\mathbb{P}\bigg{(}\frac{1}{t_{r}-1}\sum_{q=1}^{t_{r}}\left(X_{i}(q)-\bar{X}_{i}(t_{r})\right)^{2}-$
$\frac{1}{t_{r}-1}\sum_{q=1}^{t_{r}}\left(X_{*}(q)-\bar{X}_{*}(t_{r})\right)^{2}-(\sigma^{2}_{i}-\sigma^{2}_{*})>\delta_{i}\bigg{)}$
@@
$e_{n}\leq 3\log_{2}(K)\exp\left(\frac{-\left(n-K\log_{2}{K}\right)^{2}}{8n\log_{2}(K)H_{2}}\right)$
@@
$3\sum_{r=1}^{\log_{2}(K)}\exp\left(-\frac{\left(n-4i_{r}\log_{2}(K)\right)^{2}\delta_{i_{r}}^{2}}{8ni_{r}\log_{2}(K)}\right)$
@@
$\leq 3\log_{2}(K)\exp\left(\frac{-\left(n-K\log_{2}{K}\right)^{2}}{8n\log_{2}(K)H_{2}}\right).$
@@ III-B Lower bound on error for BAI
$\mathbb{P}\left(\left|\frac{1}{n}\sum_{s=1}^{n}Z_{s}-\mathbb{E}[Z]\geq\epsilon\right|\right)\leq 2\exp\left(\frac{-2{n}\epsilon^{2}}{M^{2}}\right).$
@@
$\hskip 56.9055pt\frac{1}{6}\exp\left(-3\frac{n}{H_{\sigma}(1)\sigma_{\nu}^{2}}-\frac{M}{\sqrt{2}}\sqrt{n\log(6nK)}\right)$
where $\mathbb{P}_{k}(J_{n}\neq k)$ is the probability of error under problem $G^{k}$, $H(k)=\sum_{1\leq i\leq K,i\neq k}\delta_{i}^{-2}$ measures the difficulty of identifying the best arm in problem $G^{k}$, $H_{\sigma}(1)=H(1)\delta_{k}$.
@@ IV Sub-Gaussian Distributions and Sharpe Ratio
In this section, we extend our framework from bounded distributions to sub-Gaussian distributions. Specifically, we first derive a concentration inequality for the sample variance of sub-Gaussian random variables. Building on this, we then extend our analysis to establish novel concentration bounds for the SR. It is important to note that in this work, we define the SR as $\mu/\sigma^{2}$, where $\mu$ is the mean and $\sigma^{2}$ is the variance, deviating from the classical definition of $\mu/\sigma$.
@@
Figure 1: Regret v/s time for UCB-VV, $\epsilon$-Greedy and VTS for (a) $\delta=0.1$ and (b) $\delta=0.5$ (c) UCB-VV v/s KL-UCB.
@@ IV-B UCB-Sharpe algorithm and its regret
Leveraging the above, we propose the UCB-Sharpe algorithm that selects arms with the high SR, balancing empirical performance with uncertainty in estimation. Let each arm $i\in\{1,\dots,K\}$ have unknown mean $\mu_{i}$, variance $\sigma_{i}^{2}$, and true SR $S_{i}=\mu_{i}/\sigma_{i}^{2}$. The UCB-Sharpe index at time $t$ is $\text{Index}_{i}(t)=\bar{S}_{i}(t)+\beta_{i}(t)$, where $\bar{S}_{i}(t)=\bar{\mu}_{i}/\bar{\sigma}_{i}^{2}$ is the estimated SR and $\beta_{i}(t)$ is the confidence bonus derived from the concentration inequality for the sample SR and is equal to $\beta_{i}(t)=\sqrt{\frac{1}{cs_{i}(t)}\log(4t^{2})}$. The steps are presented in Algorithm 3.
@@
6:        Set index: $\text{Z}_{i}(t)=\bar{S}_{i}(t)+\beta_{i}(t)$
@@
$\text{R}_{\texttt{SR}}(n)=\sum_{i:\Delta_{i}>0}\Delta_{i}\cdot\mathbb{E}[s_{i}(n)],$ (4)
@@
$\text{R}_{\texttt{SR}}(n)\leq\sum_{i:\Delta_{i}>0}\left(\frac{9}{c\Delta_{i}}\log n+\mathcal{O}(1)\right).$
@@
The regret scales logarithmically with $n$, similar to classical UCB algorithms, with constants depending on the sub-Gaussian parameter $v^{2}$, $\mu$ and $\sigma^{2}$.
Figure 2: Error probability $e_{n}$ of SHVV and uniform sampling algorithm for 6 experimental setups defined.